OpenStudy (anonymous):

fred had 6 3/5 quarts of oil and used 4/7 of it. how much oil did fred use?

OpenStudy (anonymous):

6 1/35

OpenStudy (anonymous):

6 3/5 = 33/5 33/5 x 4/7 =132/35